# SpiderGen:Towards Procedure Generation For Carbon Life Cycle Assessments with Generative AI

### v1

@@ Abstract
Investigating the effects of climate change and global warming caused by GHG emissions have been a primary concern worldwide. These emissions are largely contributed to by the production, use and disposal of consumer products. Thus, it is important to build tools to estimate the environmental impact of consumer goods, an essential part of which is conducting Life Cycle Assessments (LCAs). LCAs specify and account for the appropriate processes involved with the production, use, and disposal of the products. We present SpiderGen, an LLM-based workflow which integrates the taxonomy and methodology of traditional LCA with the reasoning capabilities and world knowledge of LLMs to generate the procedural information used for LCA. We additionally evaluate the output of SpiderGen using real-world LCA documents as ground-truth. We find that SpiderGen provides accurate LCA process information that is either fully correct or has minor errors, achieving an F1-Score of 62% across 10 sample data points. We observe that the remaining missed processes and hallucinated errors occur primarily due to differences in detail between LCA documents, as well as differences in the “scope” of which auxiliary processes must also be included. We also demonstrate that SpiderGen performs better than several baselines techniques, such as chain-of-thought prompting and one-shot prompting. Finally, we highlight SpiderGen’s potential to reduce the human effort and costs for estimating carbon impact, as it is able to produce LCA process information for less than \$1 USD in under 10 minutes as compared to the status quo LCA, which can cost over \$25000 USD and take up to 21-person days.
@@ 1 Introduction
Figure 2: The SpiderGen workflow utilizes LLMs and sentence transformers to create Product Category Rule process flow graphs. This workflow consists of (1) A product generation step, where sample products are sourced and (2) processes for those sample products are generated, (3) A coarse process generation step, where the sample product processes are coarsened to create generalized processes and (4) the generation of the process flow graph itself. The formation of these graphs can then be used for a variety of applications, such as carbon footprint and supply chain analysis.
@@
Life Cycle Assessments (LCAs) are commonly used to evaluate the environmental impact of a product in all phases of the product’s “life”, such as the manufacturing phase or the use phase by the end consumer. Although LCAs are essential for measuring the environmental impacts of products, the end-to-end time and monetary cost of conducting LCAs is extraordinarily high - a 2017 study shows that just determining the appropriate processes for a given product category can cost over to \$25000 USD, and can take up 21 person days to produce (tasaki_international_2017). Due to the challenges of generating the appropriate processes for an LCA, it is very difficult to scale environmental impact assessments. As a result, LCAs are only available for a very small subset of products.
A key challenge in creating an LCA is choosing which processes to include in the analysis, and creating standardized methodologies with which to evaluate specific categories of products. Traditionally, the appropriate processes for a given product category are determined by a committee of experts, and the LCA practitioner utilizes these processes and augments them for the specific product. These processes are described using a Process Flow Graph (PFG) to represent the ordering and dependencies throughout the value chain of the product. An example of a PFG for the product category “Wine” is shown in Figure 1.
@@
To address these challenges, there has been a growing interest in using machine learning techniques to automate LCA procedures to allow for easier, faster and more ubiquitous LCAs. Prior work has primarily focused on using existing process rules for specific categories of products, “bill of materials” (BOMs), and LCA guidelines, and automating other pieces of the LCA pipeline (balaji_caml_2023; balaji_flamingo_2023; balaji_parakeet_2024; zhang_deltalca_2024; sousa_approximate_2002; Wang2025). However, to our knowledge, we are the first to explore the potential of using ML to generate PFGs for LCA. Automating this key step will enable LCAs of products which have no guiding PFG available previously, such as laptops. This problem is complex, as it requires the collection of a wide range of process information, understanding the dependencies of these processes, and the ability to generalize processes to be applicable to all products of a given category.
@@
Our Approach. We propose SpiderGen  [^2] Much like a spider, SpiderGen is a workflow which generates a web of resources and weaves insights together to generate PFGs., a novel LLM-based workflow that automates the production of PFGs for a given product category. To do so, SpiderGen utilizes the world knowledge capabilities of LLMs, as well as text-embedding methods such as SBERT (reimers-2019-sentence-bert), and graph clustering algorithms such as K-Means clustering to derive upstream, core, and downstream processes for the overarching product category. SpiderGen then orders these processes to produce a Directed Acyclic Graph (DAG) representation of the end-to-end life cycle of the product category. Notably, we leverage specific ontologies and taxonomies involved in LCA production, to be able to generate relevant processes that are required in the PFG. We specifically address the challenges of generating processes that are generalizable across all products of a given product category, ordering them based on their dependencies, and ensuring that all of the processes that may be within the “scope” of a true LCA are included.
@@
Evaluation Methodology. To evaluate SpiderGen, we utilize a ground truth dataset of 66 LCA documents from EPD International (EPD). To compare the generated PFGs to the ground truth, we introduce a novel evaluation methodology (Section 4) since prior knowledge graph evaluation methods, such as those for UML diagrams and Buisness Process Flow graphs are not appropriate for our problem (fauzan_structural_2024; faisal_approach_2018; kristina_business_2024). Complete PFGs must include all the correct processes and nothing more, and must include the correct partial ordering of processes, which reflect the dependencies of each process. To determine whether SpiderGen has generated the correct processes and ordering, we provide both a qualitative and quantitative evaluation of SpiderGen. For qualitative evaluation, we outline a set of criteria that determine “correctness” and evaluate 10 datapoints using this criteria. For the quantitative evaluation, we calculate the Pointwise Mutual Information (PMI) between the ground-truth and SpiderGen’s generated PFGs.
@@
  • We find that SpiderGen provides accurate PFGs as compared to the ground truth, producing processes that are fully correct or only with minor errors 62% of the time on average. We find that the main challenges of increasing the accuracy of SpiderGen lie in producing all of the appropriate auxiliary processes that are part of the ground truth PFG, while avoiding extrapolated processes that may not be applicable to a given product category.
@@ 2 Related Work
LCA is the method used to estimate the carbon footprint of products and processes. A LCA expert starts by identifying the processes and environmental factors of the product life cycle that must be included. To do so, LCA experts rely on Process Flow Graphs (PFGs), which describe the processes for a specific product category that must be included in the LCA and their dependencies. By utilizing these graphs, experts can avoid making errors while conducting LCAs. Further, these graphs allow products of the same category to be evaluated in a consistent way, allowing for appropriate LCA comparisons. Additionally, these graphs also prevent “greenwashing”, where environmental impact information is manipulated to make products appear more sustainable than they truly are.
@@
Incorporating machine learning techniques for different parts of the LCA process has become a growing area of interest (algren2021).
Some of the earlier efforts utilize surrogate modeling to create new LCA analyses based on a training set of LCAs for a set number of products (sousa_approximate_2002; sousa_learning_nodate; sousa_product_2006).
More recently, researchers have proposed using machine learning and LLMs to extract relevant information from LCA reports and databases (goridkov_whats_2024). Balaji et al. use sentence embeddings and LLMs to match environmental impact factors from LCA databases to a given product descriptions (balaji_flamingo_2023; balaji_caml_2023; balaji_parakeet_2024). A related body of work uses information extraction methods to synthesize LCAs by utilizing templates for a given product, such as a Bill of Materials (BOM) and domain-specific databases to guide the extraction, synthesis or prediction of LCAs (zhang_deltalca_2024). In contrast to our work, these prior works do not explore the generation of PFGs for LCA.
@@
To the best of our knowledge, we are the first to explore the problem of generating process flow graphs for LCA. In contrast to prior work, our approach is zero-shot and uses off-the-shelf models to generate PFGs. We achieve this by applying the ontology and taxonomies of LCA production to create an LLM workflow for graph construction for our specific problem. Unlike fine-tuning based methods, SpiderGen is not constrained by limited training data, which allows SpiderGen to create a PFG for almost every product category.
@@ 4 SpiderGen: Our Proposed Solution
Quantity: To gain a sufficient understanding of a product category, we must ask for the right number of products to use. In early experiments, we found that asking for too few products would result in PFG information that was over-fitted to a specific niche within a product category. However, querying for too many products may make the PFG over-generalized and therefore inaccurate.
(2) Generate Sample Product Process Lists: To find common processes across a given product category, SpiderGen instructs the LLM to select diverse products that use a wide variety of materials and processes. For each product, SpiderGen queries the LLM for details about what components make up the product, and how these components are processed. We then query the LLM to find the processes involved in manufacturing and distributing the product. We prompt the LLM to list the process name, the life cycle phase it is involved in, and describe why this process is included. We use the LCA ISO standards to guide the LLM for process generation to cover various types of processes and ensure complete coverage of environmental impacts (EPD).
(3) Generate Coarse Process List: From the detailed process descriptions of a list of products, we create coarser generalized processes that can be included in a PFG. Intuitively, similar processes that appear in all product processes should be included in the PFGs, since it is likely that they are an important process to include for the entire category. Based on this intuition, we identify clusters of common processes to include in the final PFG. Finally, we utilize an LLM to summarize these clusters to create coarse process descriptions and eliminate clusters that are repeated. To generate relevant clusters of nodes, we utilize a pre-trained SBERT model to embed each process. We then use K-means clustering to provide groupings of repeated processes. To ensure that we do not cluster processes that may seem similar but are a part of different life cycle phases within the final PFG, we only cluster processes that are in the same life cycle phase. For each life cycle phase, we select the number of clusters that minimizes the Davies Bouldin score of each group life cycle phase clusters. Using these clusters, we then prompt the LLM to give a description of each of the clusters and then remove repeated descriptions. The resultant descriptions are the coarse processes used in the final process flow graph.
@@ 5 Evaluation Methodology
We evaluate the generation of PFG $G_{pc}$ on a set of 66 Product Category Rules (PCR) documents in the EPD International Database (EPD). These documents contain generalized processes for a given product category or a set of product categories for each phase of the product life cycle (upstream, core, and downstream). We manually extract the processes from these PCR documents and form a PFG with these processes. [^3] We initially attempted to use Retrieval Augmented Generation (RAG) to extract process descriptions from the PCR documents. However, due to the lack of consistency across different documents, the RAG sometimes missed smaller processes. To ensure full correctness in the ground truth, we opted to conduct this extraction manually. We provide the code for the RAG method in our GitHub repository. Each of the PCR documents have United Nations Central Product Classification (UN CPC) codes that describe the product category. For each PCR, we compile descriptions of the relevant UN CPC codes, as well as the name of the PCR. This UN CPC information is given to SpiderGen as an input for generating the PFG. We experiment with three different LLM models: OpenAI’s GPT-4o, o1-preview and o1-mini models. We used these models for our evaluation in April, May, June and July of 2025.
@@
We qualitatively analyze the results for 10 different product categories by comparing the PFGs generated by SpiderGen to the ground truth PFGs, categorizing any errors that SpiderGen makes. This procedure involves two steps: We first pair all of the nodes by similarity by pairing every node from the generated graph $G_{pc}$ to the most similar node from the ground truth graph. Every node from both graphs must be included in the matching, and many-to-many matching are allowed. If there is no appropriate pairing for a given node, we pair the node with a fake node called ”N/A”.
@@
  • If there exists a generated node that is an inferred (but not explicitly stated) sub-process of a ground-truth node, this match is labeled as “subprocess” (i.e, if there is a ground-truth node called “Manufacturing Pasta” and a generated node called “Cutting pasta dough”, this is a “subprocess”)
@@
$\text{Recall}=(c_{specific}+c_{subprocess}+c_{match})/(n_{groundtruth})$ (2)
@@
$\text{Precision}=1-(c_{wrong})/(n_{generated})$ (3)
where $c_{specific}$ is the count of ground truth nodes that are in “specific” pairings, $c_{subprocess}$ is the count of ground-truth nodes that are in “subprocess” pairings, $c_{match}$ is the count of ground truth nodes that are in “match” pairings, $n_{groundtruth}$ is the total number of ground truth nodes, $c_{wrong}$ is the count of generated nodes that are labeled as “wrong”, and $n_{generated}$ is the total number of generated nodes.
We provide a small sample of manually evaluated [^4] We initially utilized LLM-as-a-Judge techniques to rapidly conduct multiple of these evaluations. However, we observed that even with the addition of example cases, and chain-of-thought prompting, LLM-as-a-Judge tended to estimate a higher recall than a human judge. analyses, and describe themes in the erroneous processes produced by SpiderGen. This qualitative analysis for 10 product categories is given in Table 1. We observe that the majority of errors produced by SpiderGen derive from two sources: (1) misalignment of an appropriate “scope” for the processes that must be included (i.e, how many tangential or auxiliary processes should be included), and (2) extrapolation and assumptions about how a product might be used and what materials may be involved to create the product.
@@
  • Pre-processing both lists: We organize both lists such that upstream processes are listed first, core processes second, and downstream processes third. Both lists start with “raw material procurement” processes, and end with “end-of-life” processes. For ground-truth documents, direct references to other documents, links, references to other sections of the text and references to calculations have been removed. Additionally, processes marked as “optional” are marked as such. Xu et.al propose “pre-processing” text by using an LLM to rephrase the texts before calculating PMI. However, this pre-processing is not feasible for our problem as it may lead to a higher risk of introducing hallucinations that can change the meaning and structure of the ground truth text.
@@
Product Category PMI F1-Score
Railways (44) 0.02 0.37
Shower Enclosures (40) 0.02 0.66
T-Shirts, Tops (28) 0.03 0.63
Moka Coffee (22) 0.02 0.63
Dairy Products (22) 0.05 0.61
Graphite Products (20) 0.04 0.63
Detergents & Washing (19) 0.09 0.50
Woven Fabric (17) 0.05 0.84
Air Ducts (17) 0.07 0.71
Bottled Water (12) 0.03 0.59
@@
Figure 4: Comparing the Normalized PMI Scores for 66 products with varying complexity as denoted by the number of nodes in the ground truth $G_{pc}$ graphs. SpiderGen has higher normalized PMI for simpler product categories (e.g. “Dairy Products”) and lower for more complicated ones (e.g. “Railways”). We also report the qualitative scores for 10 different product categories (circled nodes) in Table 4 .

### v2

@@ Abstract
Investigating the effects of climate change and global warming caused by GHG emissions have been a key concern worldwide. These emissions are largely contributed to by the production, use and disposal of consumer products. Thus, it is important to build tools to estimate the environmental impact of consumer goods, an essential part of which is conducting Life Cycle Assessments (LCAs). LCAs specify and account for the appropriate processes involved with the production, use, and disposal of the products. We present SpiderGen, an LLM-based workflow which integrates the taxonomy and methodology of traditional LCA with the reasoning capabilities and world knowledge of LLMs to generate graphical representations of the key procedural information used for LCA, known as Product Category Rules Process Flow Graphs (PCR PFGs). We additionally evaluate the output of SpiderGen by comparing it with 65 real-world LCA documents. We find that SpiderGen provides accurate LCA process information that is either fully correct or has minor errors, achieving an F1-Score of 65% across 10 sample data points, as compared to 53% using a one-shot prompting method. We observe that the remaining errors occur primarily due to differences in detail between LCA documents, as well as differences in the “scope” of which auxiliary processes must also be included. We also demonstrate that SpiderGen performs better than several baselines techniques, such as chain-of-thought prompting and one-shot prompting. Finally, we highlight SpiderGen’s potential to reduce the human effort and costs for estimating carbon impact, as it is able to produce LCA process information for less than \$1 USD in under 10 minutes as compared to the status quo LCA, which can cost over \$25000 USD and take up to 21-person days.
@@ 1 Introduction
Figure 2: The SpiderGen workflow utilizes LLMs and sentence transformers to create Product Category Rule Process Flow Graphs (PCR PFGs). This workflow consists of (1) A product generation step, where sample products are sourced and (2) processes for those sample products are generated, (3) A coarse process generation step, where the sample product processes are coarsened to create generalized processes and (4) the generation of the process flow graph itself. The formation of these graphs can then be used for a variety of applications, such as generating carbon footprints and and conducting supply chain analysis.
@@
Life Cycle Assessments (LCAs) are commonly used to evaluate the environmental impact of a product in all phases of the product’s “life”, such as the manufacturing phase or the use phase by the end consumer. Although LCAs are essential for measuring the environmental impacts of products, the end-to-end time and monetary cost of conducting LCAs is extraordinarily high - a 2017 study shows that just determining the appropriate processes for a given product category can cost over \$25000 USD, and can take up 21 person days to produce (tasaki_international_2017). Due to the challenges of generating the appropriate processes for an LCA, it is very difficult to scale environmental impact assessments. As a result, LCAs are available for a very small subset of products.
@@
A key challenge in creating an LCA is choosing which processes to include in the analysis, and creating standardized methodologies with which to evaluate specific categories of products. Traditionally, the appropriate processes for a given product category are determined by a committee of experts, and the LCA practitioner utilizes these processes and augments them for the specific product. These processes are described using a Process Flow Graph (PFG) to represent the ordering and dependencies throughout the value chain of each product. An example of a PFG for the product category “Wine” is shown in Figure 1. To address this challenge, there has been a growing interest in using machine learning techniques to automate LCA procedures to allow for easier, faster and more ubiquitous LCAs. Prior work has primarily focused on using existing process rules for specific categories of products, “bill of materials” (BOMs), and LCA guidelines, and automating other parts of the LCA pipeline (balaji_caml_2023; balaji_flamingo_2023; balaji_parakeet_2024; zhang_deltalca_2024; sousa_approximate_2002; Wang2025). However, to our knowledge, we are the first to explore the potential of using ML to generate PFGs for LCA. Automating this key step will enable LCAs of products which have no guiding PFG available previously, such as toothbrushes. This problem is complex, as it requires the collection of a wide range of process information, understanding the dependencies of these processes, and the ability to generalize processes to be applicable to all products of a given category.
@@
Our Approach. We propose SpiderGen  [^2] Much like a spider, SpiderGen is a workflow which generates a web of resources and weaves insights together to generate PFGs., a novel LLM-based workflow that automates the production of PFGs for a given product category. To do so, SpiderGen utilizes the world knowledge capabilities of LLMs, as well as text-embedding methods such as SBERT (reimers-2019-sentence-bert), and graph clustering algorithms such as K-Means clustering to derive upstream, core, and downstream processes for the overarching product category. SpiderGen then orders these processes to produce a Directed Acyclic Graph (DAG) representation of the end-to-end life cycle of the product category. Notably, we leverage specific ontologies and taxonomies involved in LCA production (EPD; epa_pcr_us_2024), to be able to generate relevant processes that are required in the PFG. We specifically address the challenges of generating processes that are generalizable across all products of a given product category, ordering them based on their dependencies, and ensuring that all of the processes that may be within the “scope” of a true LCA are included.
@@
Evaluation Methodology. To evaluate SpiderGen, we utilize a ground truth dataset of 65 LCA documents from EPD International (EPD). To compare the generated PFGs to the ground truth, we introduce a novel evaluation methodology (Section 4) since prior knowledge graph evaluation methods, such as those for UML diagrams and Business Process Flow graphs are not appropriate for our problem (fauzan_structural_2024; faisal_approach_2018; kristina_business_2024). Complete PFGs must include all the correct processes and nothing more, and must include the correct partial ordering of processes, which reflect the dependencies of each process. To determine whether SpiderGen has generated the correct processes and ordering, we provide both a qualitative and quantitative evaluation of SpiderGen. For the quantitative evaluation, we calculate the Pointwise Mutual Information (PMI) (xu_benchmarking_2025) between the ground-truth and SpiderGen’s generated PFGs for all 65 LCA documents. For qualitative evaluation, we outline a set of criteria that determine “correctness” and evaluate 10 of the 65 generated PFGs using this criteria.
@@
  • We show that SpiderGen accurately produces PFGs relative to the ground truth, yielding an average F1-score of 65%, as compared to an average F1-score of 53% using a one-shot prompting method. We find that the main challenges of increasing the accuracy of SpiderGen lie in producing all of the appropriate auxiliary processes that are part of the ground truth PFG, while avoiding extrapolated processes that may not be applicable to a given product category.
@@ 2 Related Work
LCA is the method used to estimate the carbon footprint of products and processes. A LCA expert starts by identifying the processes and environmental factors of the product life cycle that must be included. To do so, LCA experts rely on Process Flow Graphs (PFGs), which describe the processes for a specific product category that must be included in the LCA and their dependencies. By utilizing these graphs, experts can avoid making errors while conducting LCAs. Further, these graphs allow products of the same category to be evaluated in a consistent way, allowing for appropriate LCA comparisons. Additionally, these graphs also prevent “greenwashing”, where environmental impact information is manipulated to make products appear more sustainable than they truly are (brandao_life_2024).
@@
Incorporating machine learning techniques for different parts of the LCA process has become a growing area of interest (algren2021). Some of the earlier efforts utilize surrogate modeling to create new LCA analyses based on a training set of LCAs for a set number of products (sousa_approximate_2002; sousa_learning_nodate; sousa_product_2006). More recently, researchers have proposed using machine learning and LLMs to extract relevant information from LCA reports and databases (goridkov_whats_2024). Balaji et al. use sentence embeddings and LLMs to match environmental impact factors from LCA databases to a given product descriptions (balaji_flamingo_2023; balaji_caml_2023; balaji_parakeet_2024). A related body of work uses information extraction methods to synthesize LCAs by utilizing templates for a given product, such as a Bill of Materials (BOM) and domain-specific databases to guide the extraction, synthesis or prediction of LCAs (zhang_deltalca_2024). In contrast to our work, these prior works do not explore the generation of PFGs for LCA.
@@
To the best of our knowledge, we are the first to explore the problem of generating process flow graphs for LCA. In contrast to prior work, our approach is zero-shot and uses off-the-shelf models to generate PFGs. We achieve this by applying the ontology and taxonomies of LCA production to create an LLM workflow for graph construction for our specific problem. Unlike fine-tuning based methods, SpiderGen is not constrained by limited training data, which allowing SpiderGen to create a PFG for any product category.
@@ 4 SpiderGen: Our Proposed Solution
Quantity: To gain a sufficient understanding of a product category, we must ask for the right number of products to use. In early experiments, we found that asking for too few products would result in PFG information that was over-fitted to a specific niche within a product category. However, querying for too many products may make the PFG over-generalized and therefore inaccurate.(2) Generate Sample Product Process Lists: To find common processes across a given product category, SpiderGen instructs an LLM to select diverse products that use a wide variety of materials and processes. For each product, SpiderGen queries an LLM for details about what components make up the product, and how these components are processed. We then query an LLM to find the processes involved in manufacturing and distributing the product. We prompt an LLM to list the process name, the life cycle phase it is involved in, and describe why this process is included. We use the LCA ISO standards to guide the LLM for process generation to cover various types of processes and ensure complete coverage of environmental impacts (EPD).
(3) Generate Coarse Process List: From the detailed process descriptions for a list of products, we create coarser generalized processes that can be included in a PFG. Intuitively, similar processes that appear in all product processes should be included in the PFGs, since it is likely that they are an important process to include for the entire category. Based on this intuition, we identify clusters of common processes to include in the final PFG. Finally, we utilize an LLM to summarize these clusters to create coarse process descriptions and eliminate clusters that are repeated. To generate relevant clusters of nodes, we utilize a pre-trained SBERT model to embed each process. We then use K-means clustering to provide groupings of repeated processes. To ensure that we do not cluster processes that may seem similar but are a part of different life cycle phases within the final PFG, we only cluster processes that are in the same life cycle phase. For each life cycle phase, we select the number of clusters that minimizes the Davies Bouldin score of each group life cycle phase clusters. Using these clusters, we then prompt the LLM to give a description of each of the clusters and then remove repetitions. The resultant descriptions are the coarse processes used in the final process flow graph.
@@ 5 Evaluation Methodology
We evaluate the generation of PFG $G_{pc}$ on a set of 65 Product Category Rules (PCR) documents in the EPD International Database (EPD). These documents contain generalized processes for a given product category or a set of product categories for each phase of the product life cycle (upstream, core, and downstream). We manually extract the processes from these PCR documents and form a PFG with these processes. Each of the PCR documents have United Nations Central Product Classification (UN CPC) codes that describe the product category. For each PCR, we compile descriptions of the relevant UN CPC codes, as well as the name of the PCR. This UN CPC information is given to SpiderGen as an input for generating the PFG. We experiment with three different LLM models: OpenAI’s GPT-4o, o1-preview and o1-mini models. We used these models for our evaluation from April to July 2025.
@@
We qualitatively analyze the results for 10 different product categories by comparing the PFGs generated by SpiderGen to the ground truth PFGs, categorizing any errors that SpiderGen makes. This procedure involves two steps: We first pair all of the nodes by similarity by pairing every node from the generated graph $G_{pc}$ to the most similar node from the ground truth graph. Every node from both graphs must be included in the matching, and many-to-many matching are allowed. If there is no appropriate pairing for a given node, we pair the node with a fake node called “N/A”.
@@
  • If there exists a generated node that is an inferred (but not explicitly stated) sub-process of a ground-truth node, this match is labeled as “subprocess” (i.e, if there is a ground-truth node called “Manufacturing Pasta” and a generated node called “Cutting pasta dough”, this is a “subprocess”).
@@
$\text{Recall}=(n_{specific}+n_{subprocess}+n_{match})/(n_{groundtruth})$ (2)
@@
$\text{Precision}=1-(n_{wrong})/(n_{generated})$ (3)
where $n_{specific}$ is the number of ground-truth nodes that are in “specific” pairings, $n_{subprocess}$ is the number of ground-truth nodes that are in “subprocess” pairings, $n_{match}$ is the number of ground-truth nodes that are in “match” pairings, $n_{groundtruth}$ is the total number of ground truth nodes, $n_{wrong}$ is the number of generated nodes that are labeled as “wrong”, and $n_{generated}$ is the total number of generated nodes.We provide a set of ten manually evaluated [^3] We initially utilized LLM-as-a-Judge techniques to rapidly conduct multiple of these evaluations. However, we observed that even with the addition of example cases, and chain-of-thought prompting, LLM-as-a-Judge tended to estimate a higher recall than a human judge. Additional details of this analysis can be found in the full version of the paper (synergylabs_spidergen) analyses, and describe themes in the erroneous processes produced by SpiderGen in Table 1. We observe that the majority of errors produced by SpiderGen come from two sources: (1) misalignment of an appropriate “scope” for the processes that must be included (i.e, how many tangential or auxiliary processes should be included), and (2) extrapolation and assumptions about how a product might be used and what materials may be involved to create the product.
@@
  • Pre-processing both lists: We organize both lists such that upstream processes are listed first, core processes second, and downstream processes third. Both lists start with “raw material procurement” processes, and end with “end-of-life” processes. For ground-truth documents, direct references to other documents, links, references to other sections of the text and references to calculations have been removed. Additionally, “optional” processes are marked as such. Xu et.al propose “pre-processing” text by using an LLM to rephrase the texts before calculating PMI (xu_benchmarking_2025). However, this pre-processing is not feasible for our problem as it may increase the risk of introducing hallucinations that change the meaning and structure of the ground-truth text.
@@
Category ID Product Category PMI F1-Score
$c_{1}$ Railways (44) 0.02 0.4
$c_{2}$ Shower Enclosures (40) 0.02 0.73
$c_{3}$ T-Shirts, Tops (28) 0.03 0.77
$c_{4}$ Moka Coffee (22) 0.02 0.66
$c_{5}$ Dairy Products (22) 0.05 0.63
$c_{6}$ Graphite Products (20) 0.04 0.61
$c_{7}$ Detergents & Washing (19) 0.09 0.59
$c_{8}$ Woven Fabric (17) 0.05 0.84
$c_{9}$ Air Ducts (17) 0.07 0.65
$c_{10}$ Bottled Water (12) 0.03 0.59
@@
Figure 4: Comparing the Normalized PMI Scores for 65 products with varying complexity as denoted by the number of nodes in the ground truth $G_{pc}$ graphs. SpiderGen has higher normalized PMI for simpler product categories (e.g. “Dairy Products”) and lower for more complicated ones (e.g. “Railways”). We also report the qualitative scores for 10 different product categories (circled nodes) in Table 1, and include labels for some of the product categories in Table 1 in this figure.